OpenStudy (anonymous):

Solve the equation. 2(x+4)=30 A. 7 B. 11 C. 15 D. 19

sam (.sam.):

Multiply 2 into brackets

sam (.sam.):

2x+8=30

sam (.sam.):

what you dont understand from here?

OpenStudy (anonymous):

I wanted to know how do you solve for x ?

OpenStudy (king):

2x+8=30 2x=22 x=11

sam (.sam.):

2(x+4)=30 Divide each term in the equation by 2. (2(x+4))/(2)=(30)/(2) Remove the common factors that were cancelled out. (x+4)=(30)/(2) Remove the parentheses around the expression x+4. x+4=(30)/(2) Remove the common factors that were cancelled out. x+4=15 Since 4 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 4 from both sides. x=-4+15 x=11